### v1

# Advancing Neural Network Performance through Emergence-Promoting Initialization Scheme

Abstract
We introduce a novel yet straightforward neural network initialization scheme that modifies conventional methods like Xavier and Kaiming initialization. Inspired by the concept of emergence and leveraging the emergence measures proposed by Li (2023), our method adjusts the layer-wise weight scaling factors to achieve higher emergence values. This enhancement is easy to implement, requiring no additional optimization steps for initialization compared to GradInit. We evaluate our approach across various architectures, including MLP and convolutional architectures for image recognition, and transformers for machine translation. We demonstrate substantial improvements in both model accuracy and training speed, with and without batch normalization. The simplicity, theoretical innovation, and demonstrable empirical advantages of our method make it a potent enhancement to neural network initialization practices. These results suggest a promising direction for leveraging emergence to improve neural network training methodologies. Code is available at: https://github.com/johnnyjingzeli/EmergenceInit.
1 Introduction
Emergence in machine learning refers to the phenomenon where complex behaviors and properties arise from the interactions of simpler elements within a system. This concept, deeply rooted in the study of complex systems, has gained increasing attention in the context of neural networks. The emergent properties are not merely the sum of individual components but rather the result of intricate interactions among them, leading to sophisticated and often unexpected capabilities. These emergent behaviors are crucial for enabling neural networks to perform complex tasks such as image recognition, natural language processing, and strategic game playing.
The concept of emergence is particularly relevant in neural networks, where the interplay of numerous neurons can give rise to high-level abstractions and decision-making capabilities. Li (2023) provides a framework to quantify emergence in neural networks through the measure of paths from inactive to active nodes. This metric offers a novel lens through which we can understand and enhance the emergent properties of neural networks, potentially leading to more efficient and effective models.
In this paper, we introduce a new initialization scheme for neural networks that leverages the concept of emergence. Our method adjusts the layer-wise variance parameters to achieve higher emergence values compared to traditional methods like Xavier and Kaiming initialization. This approach is particularly appealing because it is straightforward to implement, requiring only minor modifications to existing initialization techniques without necessitating additional optimization or training steps.
Our initialization scheme is grounded in the idea that by enhancing the emergent properties of neural networks from the beginning, we can facilitate better feature differentiation and integration. This, in turn, can lead to improved performance across various tasks and architectures. We evaluate our method on both convolutional neural networks (CNNs) for image recognition and transformer architectures for machine translation, demonstrating significant improvements in model accuracy and convergence speed.
Emergence in neural networks can be seen as a manifestation of the broader principle that simple rules or interactions can lead to complex and adaptive behaviors. This principle has been explored in various contexts within machine learning, from the self-organizing properties of unsupervised learning algorithms to the adaptive behaviors of reinforcement learning agents. The work by Li (2023) extends this understanding by providing a concrete measure of emergence that can be directly applied to the design and initialization of neural networks.
Our experiments on standard datasets such as ImageNet for image recognition and the WMT news translation task for machine translation underscore the practical benefits of our initialization scheme. The results highlight not only improvements in final model accuracy but also faster convergence during training, suggesting that our method effectively primes the network for learning by optimizing its emergent properties from the outset.
The simplicity and effectiveness of our approach make it a compelling addition to the toolkit of neural network initialization methods. By focusing on enhancing emergent properties, our scheme offers a new perspective on how initialization can impact the learning dynamics and ultimate performance of neural networks. This paper contributes to the growing body of research that seeks to understand and harness the power of emergence in machine learning, paving the way for more robust and capable models.
In summary, our proposed initialization scheme represents a novel and practical approach to improving neural network performance by leveraging the concept of emergence. By adjusting layer-wise variance parameters to enhance emergent properties, we provide a simple yet powerful modification to existing initialization methods. The empirical results presented in this paper demonstrate the significant advantages of our approach, offering a promising direction for future research and applications in machine learning.
2 Related Work
The initialization of neural networks has been a critical area of research, influencing the stability and speed of training, as well as the ultimate performance of the models. Traditional initialization schemes such as Xavier[12] and Kaiming [13] initialization have laid the foundation for effectively training deep networks by mitigating issues related to vanishing and exploding gradients. Xavier initialization aims to keep the scale of the gradients approximately the same in all layers, while Kaiming initialization, specifically designed for ReLU activations, helps to maintain the variance of activations throughout the layers. Both methods have proven to be fundamental in training deep networks but do not explicitly account for emergent properties within the networks.
Recent research has explored more sophisticated initialization strategies that leverage the structural and statistical properties of neural networks. For instance, Saxe et al. (2013) studied the dynamics of signal propagation in deep networks, highlighting the importance of properly scaling the initial weights to ensure efficient training. Additionally, Mishkin and Matas (2015) proposed a layer-sequential unit-variance (LSUV) initialization that iteratively adjusts the weights to achieve unit variance across all layers, further improving convergence.
The concept of emergence, where complex behaviors arise from simple interactions within a system, has also been examined in the context of neural networks. Emergent properties have been shown to play a crucial role in the development of robust and adaptive models. Research by Olah et al. (2020) has illustrated how higher-level features and behaviors emerge in deep networks as a result of training on large datasets. This phenomenon underscores the potential for leveraging emergent properties to enhance network performance.
(Li, 2023) introduced a novel framework for measuring emergence in neural networks, focusing on the number of paths from inactive to active nodes as a key metric. This work provides a theoretical basis for understanding how emergent behaviors can be quantified and optimized, offering new insights into the design and initialization of neural networks.
Our proposed initialization scheme builds on these foundational studies by explicitly targeting the enhancement of emergent properties from the outset. By adjusting layer-wise variance parameters to achieve higher emergence values, our method represents a simple yet effective modification to traditional initialization techniques. This approach aligns with the growing body of research that seeks to integrate principles from complex systems and emergence theory into machine learning practices.
Furthermore, our work is related to studies that explore the role of initialization in specific architectures such as transformers. Vaswani et al. (2017) introduced the transformer model, which has become a cornerstone in natural language processing due to its ability to capture long-range dependencies through self-attention mechanisms. Recent research has continued to refine transformer architectures, with improvements in initialization playing a crucial role in achieving state-of-the-art performance (Liu et al., 2020).
In summary, our work contributes to the ongoing exploration of initialization methods by introducing a scheme that enhances the emergent properties of neural networks. By drawing on the principles of emergence and leveraging recent advancements in the field, we provide a novel perspective on how initialization can impact the learning dynamics and performance of deep models across various architectures and tasks.
3 Method
Emergence in machine learning refers to the phenomenon where complex, higher-level behaviors and structures arise from the interactions of simpler elements within a system. This concept, rooted in various scientific disciplines, has been increasingly recognized for its potential to enhance our understanding and development of neural networks. In our formulation, emergence is quantified through a graph-theoretical framework, which examines the network topology and the interactions between nodes to provide a clear and structural metric for emergent behavior.
A structural definition of emergence offers several key benefits. First, it allows for a more precise and quantifiable understanding of how complex behaviors develop in neural networks. This is particularly useful for designing and optimizing network architectures, as it provides clear metrics that can be targeted during the training and initialization phases. Second, a structural approach facilitates better theoretical insights, enabling researchers to draw more concrete connections between network topology and emergent properties. This can lead to more effective and efficient network designs, ultimately improving performance across various tasks.
Emergence in this context is inherently multiscale. It involves interactions across different scales of the network, where $G$ represents one scale and $H$ represents a higher scale. Emergence appears only when viewed from this multiscale perspective, as it captures the complexity arising from the network’s hierarchical structure.
In our graph-theoretical framework, the emergence value $E$ of a neural network is defined based on the number of paths from nodes at scale $G$ to nodes at scale $H$. This definition captures the essence of multiscale interactions within the network. The more paths that exist between these scales, the greater the degree of emergence.
Mathematically, this can be expressed as:
$\text{Emergence}(G,H)$ $=\sum_{x\in G\setminus H}\#\text{paths in $G$ from $N_{H}(x)$ to $H$}.$ (1)
where:
  • $G$ is the set of nodes at the lower scale,
  • $H$ is the set of nodes at the higher scale,
This equation captures the emergent behavior by accounting for the connections and interactions between different scales within the network. The higher the value of $E$, the more interconnected the network is across scales, thereby increasing the likelihood of complex behaviors and traits emerging from the network.
Intuitively, a system with a higher value of $E$ has more extensive and interconnected pathways through which information can propagate across different scales. This interconnectedness facilitates the development of intricate patterns and features within the network, enabling it to capture and represent more complex relationships in the data. As a result, networks with higher emergence values are better equipped to learn and generalize from diverse and intricate datasets, leading to improved performance across various machine learning tasks.
Our approach leverages this definition to modify the initialization process of neural networks, aiming to enhance their emergent properties from the outset. By doing so, we can achieve significant improvements in network performance, as demonstrated in our experimental results.
To understand emergence in the context of machine learning, when a model has stronger emergence traits, this means that the model is easy to learn any or certain downstream tasks. From a loss function perspective, a model with a stronger emergence should be closer to the global minimum, or the learning should be fast. We will show in our numerical experiments, that schemes with stronger emergence will indeed have faster learning in the initial epochs.
To adapt the definition of emergence in our context, we should give a clear definition of $G$ and $H$ in the mathematical definition above. Since emergence here evaluates the potential/ ability for emergent traits when we observe system $G$ from a higher level $H$, here we want $G$ to represent the model itself, and $H$ to be some certain features of the model. In the paper, we adopt the setting that $H$ is the nodes in $G$ that are still active in the training process, where the criteria for active nodes is the set of nodes whose activation is greater than a threshold set close to 00. This sorted out the nodes that are not actively participating in the computational process. The set of active nodes thus in a sense represent the learning task, thus we can tie emergence with the performance of the network in a learning process. This fits in our framework of emergence, where part of the system is being neglected after the learning process, thus the learning process represents the $\Phi$ where partial observation is carried out, and the properties of $H$ represents the emergent abilities of the network.
For a feedforward network, with $N$ layers, and $n_{i}$ nodes for each layer, and $a_{i}$ the number of active nodes for each layer, emergence is computed as follows:
$E$ $=\sum_{i=1}^{N-1}\sum_{j>i}^{N}\#\text{paths from inactive nodes in layer }i\text{to active nods in layer }j$ (2)
$=\sum_{i=1}^{N-1}\sum_{j>i}^{N}(n_{i}-a_{i})n_{i+1}\cdots n_{j-1}a_{j}$
$=\sum_{i=1}^{N-1}\sum_{j>i}^{N}(n_{i}-a_{i})a_{j}\prod_{k=i+1}^{j-1}n_{k}$
when the network architecture is fixed, which means, when $L$ and $n_{i},i=1,\ldots,L$ are fixed, emergence is only a function of the number of active nodes in each layer,
$E:=E(a_{1},\ldots,a_{N}).$ (3)
And the number of active nodes at initialization, is impacted by the weights. With a criteria for active nodes, for example, those nodes whose activation is greater than a threshold as adopted in this paper, we can thus establish initialization scheme that has stronger emergence.
Lemma: Emergence function $E(a_{1},\ldots,a_{N})$ increases when $a_{1}\ldots a_{i}$ gets smaller and $a_{i}\ldots a_{N}$ gets larger, where $i$ is the largest integer such that
$\sum_{k=1}^{N}a_{k}-a_{i}-\sum_{k=1}^{i-1}n_{i}>0.$ (4)
The proof is simple: when $a_{i}$ decrease by $1$, then the number of paths from layer $i$ to the dead nodes in latter layers will increase by $a_{i+1}+\cdots+a_{N}$, and the number of paths from earlier layers to layer $i$ will decrease by $(n_{1}-a_{1})+\cdots+(n_{i-1}-a_{i-1})$.
So the net increase of path will be
$\Delta$ $=a_{i+1}+\cdots+a_{N}-[(n_{1}-a_{1})+\cdots+(n_{i-1}-a_{i-1})]$ $=\sum_{k=1}^{N}a_{k}-a_{i}-\sum_{k=1}^{i-1}n_{i}$ (5)
For the usual neural network structures, where for each layer most nodes are active, and later layers are smaller, the balance usually occurs when $i\approx N/2$.
Let us consider now why the network with the configuration above has stronger emergence: for example, when the network is doing an image recognition task, the nodes in the later half of the layers are making important decisions on which category the image belongs to, so it needs to be more alive/ subject to more sensitive weight changes. However, for the nodes in the initial half of the layers, they are subject to greater weight changes, and they could be turned off to represent some global features of the image. Hence they could be more inactive/ subject to more drastic weight changes.
This idea also agrees with the fine-tuning idea: typically, the initial layers (closer to the input) have smaller learning rates, while the later layers (closer to the output) have larger learning rates. This strategy is based on the idea that the initial layers capture more generic features that are less likely to change significantly, whereas the later layers capture more task-specific features that require more significant adjustments. When the model has stronger emergence according to our theory, it is more likely to learn the specific tasks faster in a fine tuning process.
Now we aim at proposing a network initialization architecture with stronger emergence. To do so, we decrease the weight magnitude in the first half of the layers and increase the weight magnitude in the second half of the layers.
  • Decrease the activity of nodes in the first half layers
  • Increase the activity of nodes in the second half layers
in order to achieve this, we design the following initialization scheme:
  • Decrease the magnitude of weights in the first half layers by dividing a factor
  • Increase the magnitude of weights in the second half layers by multiplying a factor
because in general, a larger weight magnitude can lead to higher activation thus increase the number of active nodes.
We also want to consider the stability of the network. The existing initialization schemes were usually designed such that the activation and gradients are stable across the layers. For example, in Xavier initialization, it is shown that
$n_{i}Var(W_{i})=1$ (6)
$n_{i+1}Var(W_{i})=1$ (7)
promotes stability of activation and gradients. However, when we increase the variance in the inital layers and decrease the variance in the later layers, we introduce instability to the flow of activation and gradients across the layers. To reduce the effect of instability on the performance of initialization scheme, we make the increase of variance across the layers to be smooth, so as to reduce the instability. For example, we initialize the weight matrices in the following way:
We first initialize the network weights $\{W_{i}\}$ following some standard initialization scheme which preserves stability, for example, Xavier or Kaiming He Initialization. Then we do the following scaling to the weights:
$\tilde{W}_{-n}$ $=W_{n}/\alpha^{n}$ (8)
$\tilde{W}_{-(n-1)}$ $=W_{-(n-1)}/\alpha^{n-1}$
$\vdots$
$\tilde{W}_{0}$ $=W_{0}$
$\tilde{W}_{1}$ $=W_{1}*\alpha$
$\vdots$
$\tilde{W}_{n-1}$ $=W_{n-1}*\alpha^{n-1}$
$\tilde{W}_{n}$ $=W_{n}*\alpha^{n}$
In our experiments, we see (Figure 1) that this initialization is indeed leading to a better performance. In particular, we show the correlation between emergence and performance. Based on our theory, we have an increase in emergence, even when only the magnitude of weights of the first half layers decreases, or the magnitude of weights of the second half increases.
We also note that, such choice of layer magnitudes is mimicking the "domino effect", as illustrated in the figure below. The increase of energy level for each piece can set off a cascade effect.
Figure 1: The increase of weight magnitude through the layers is mimicking how the energy level through the pieces of a domino is increasing.
Figure 2: Comparison of training loss and emergence of Xavier, Kaiming and our initialization schemes.
Emergence value: Xavier: $5.03e^{8}$, Kaiming: $5.99e^{8}$, Ours: $10.87e^{8}$
We now study how to choose the scaling factor $\alpha$ properly. As shown in our numerical experiments, we can see that as $\alpha$ increase, the performance first increases then decrease, and the decrease part is likely to be caused by the instability inherent to the initialization scheme. In order to determine a factor $\alpha$ that is appropriate, we want to limit the emergence of the model to a range.
Choice of optimal $\alpha$: From the mathematical equation of emergence, we can see that given a model, the maximum amount of emergence is determined by the parameters of the equations, $N$ and $n_{i}$, which are the number of layers in the network and the size of each layer. Emergence increases in $O(n_{i})$ and $O(N^{2})$. As a result, emergence is more sensitive to $\alpha$ when the network has more layers. So we allow larger $\alpha$ when the network is shallow and smaller $\alpha$ when the network is deep. Empirically, under the learning rate $lr=0.001$, $\alpha=2$ is a good choice for usual architectures. For the two layer MLP, for example in transformers, $\alpha$ can be as large as $10$, which for deeper MLP, $n>5$, then smaller $\alpha$ should be considered.
Here our motivation is simply to bound the emergence value. We should also bear in mind that stability is also a very important issue for an initialization scheme to behave well. We encourage people to give more rigorous analysis on emergence and stability so as to strike a more optimal balance between these two.
On learning rates
Empirically, the scaling factor $\alpha$ is also dependent on learning rate. When the learning rate is faster, more stability is usually required and hence we need smaller emergence. There has been some theoretical results on the optimal learning rate, for example, Hettinger et al. (2020) suggested that the optimal learning rate should be inversely proportional to the gradient magnitude at initialization:
$\eta=\frac{c}{\|\nabla L\|}$ (9)
where $\eta$ is the optimal learning rate, $\|\nabla L\|$ is the magnitude of the gradient of the loss function $L$ with respect to the network parameters at initialization, and $c$ is a constant.
So when given the learning rate, we should choose $\alpha$ such that the resulting gradient magnitude at initialization is inversely proportional.
If we assume that by introducing our scheme we have $C(\alpha)^{N}$ times increase to the initial gradient.Then under the learning rate $\eta$ we have
$\frac{\eta}{\eta_{0}}=\frac{\|\nabla L_{\alpha_{0}}\|}{\|\nabla L_{\alpha}\|}=\frac{\alpha_{0}^{N}}{\alpha^{N}}$ (10)
hence we have
$\alpha=\alpha_{0}\Big{[}\frac{\eta_{0}}{\eta}\Big{]}^{1/N}.$ (11)
For a two layer network, if we choose $\alpha_{0}=2$ for learning rate $\eta_{0}=1e-3$, then for a different learning rate $\eta=1e-4$ we should choose $\alpha=6.32$.
Since based on our scheme, the initial gradient varies across layers, the layer-wise learning rates configuration should also be considered a good choice. Specifically, the learning rate for each layer, denoted by $\eta_{l}$, should be proportional to the inverse of the square root of the expected squared gradient norm at initialization. Mathematically, this can be expressed as:
$\eta_{l}\propto\frac{1}{\sqrt{E\left[\left\|\nabla L(\mathbf{x}_{l}^{0})\right\|^{2}\right]}}$
where $\eta_{l}$ is the learning rate for layer $l$, and $E\left[\left\|\nabla L(\mathbf{x}_{l}^{0})\right\|^{2}\right]$ is the expected squared gradient norm of the loss $L$ with respect to that layer’s inputs $\mathbf{x}_{l}^{0}$ at initialization. This approach aims to optimize the learning process by adjusting the learning rates according to the variability and scale of the gradients encountered in different layers of the neural network.
Other architectures and block wise initialization
Note that the mathematical equation (2) only works for MLP. For convolutional layers, we have the following:
$E$ $=\sum_{i=1}^{N-1}\sum_{j>i}^{N}(n_{i}-a_{i})a_{j}\prod_{k=i+1}^{j-1}m_{k}$ (12)
where $m_{k}$ is the number of filters in layer $K$. The analysis largely follows.
In most convolutional architectures and transformers, there are MLP blocks presented. While we can only do our intilization to the MLP blocks and see an improvement, we can also consider applying scheme (8) to the convolutional layers. Given this formula for emergence, we can increase the global emergence, but also increase the local emergence by applying scheme (8) to some of the layer blocks. We will discuss this in the next section.
4 experiments
We evaluate our initialization scheme on benchmark datasets for image classification and machine translation tasks. For image classification, five different architectures are evaluated for CIFAR10 [37]. For machine translation, we use our initialization scheme to find good initializations for a Post-LN Transformer without any change to its original architecture on IWSLT-14 De-En [39].
We conduct our experiments in PyTorch. We use the fairseq library for machine translation [40]. All the experiments on CIFAR-10 and IWSLT-14 DE-EN can run with one single NVIDIA A100 GPU with 13GB of RAM. Our initialization scheme first initializes the weights using Kaiming initialization for all the Conv and MLP layers for image classification. For machine translation, we use the default Xavier initialization [12]. Base on the discussion in the previous section, we choose the scale factors $\alpha=2$ with out batch normalization and $\alpha_{i}=5$ with batch normalization.
On CIFAR-10, we focus on MLP and the feedforward VGG net with and without BN layers. Since ResNet has recursive network structure, we leave it to a future work to establish the emergence formula on it. For MLP, we use a simple MLP architecture with $3$ hidden layers. For VGG net, we use VGG-19 and our initialization scheme is compared with four different methods/settings: 1) Kaiming Initialization [13]; 2) First train the network for one epoch with a constant learning rate equal to the starting learning rate, labelled as “+1 epoch (Const. LR)" in Table 3; 3) First train the network for one epoch with a linear warmup learning rate, labbeled as “+1 epoch (Warmup)" in Table 3; 4) MetaInit [27]. The data is from GradInit[8]. On CIFAR-10, we train networks with a batch size of 128, and a constant learning rate of $0.001$.
From our experiments we can see BN does stabilize VGG-19 and allows training with stronger emergence (larger value of $\alpha$).
(a) w/o batch normalization,$\alpha=2$
(b) w/ batch normazalization, $\alpha=10$
Figure 3: Training loss and test accuracy of MLP on CIFAR-10.
Table 1: First epoch (Acc1) for models on CIFAR-10, $\alpha=2$ for VGG-19 without BN, and $\alpha=5$ for VGG-19 with BN.
Model VGG-19 VGG-19 ResNet-110 ResNet-110
w/o BN w/ BN w/o BN w/ BN
Kaiming $Acc_{1}$ $29.1\pm 1.5$ $12.6\pm 0.6$ $16.1\pm 2.1$ $23.2\pm 0.9$
+1 epoch (Const. LR) $Acc_{1}$ $37.2\pm 1.1$ $19.6\pm 4.0$ $21.0\pm 3.8$ $32.5\pm 3.8$
+1 epoch (Warmup) $Acc_{1}$ $37.4\pm 1.2$ $53.5\pm 2.9$ $19.8\pm 0.5$ $48.7\pm 1.1$
MetaInit $Acc_{1}$ $30.5\pm 0.9$ $35.1\pm 0.6$ $14.6\pm 2.2$ $29.0\pm 1.5$
GradInit $Acc_{1}$ $29.3\pm 0.6$ $47.8\pm 1.8$ $36.2\pm 0.8$ $38.2\pm 0.9$
Ours $Acc_{1}$ $46.2\pm 0.6$ $52.4\pm 1.0$ $45.3\pm 2.0$ $48.0\pm 1.5$
Table 2: Evaluate the transformer model for 80 epochs
Method $\text{BLEU}_{1}$ $\text{BLEU}_{best}$
Xavier - 34.85
T-Fixup 3.96 34.78
Ours 4.8 35.13
We can see the since batch normalization promotes a good stability, we are free to choose larger $\alpha$.
IWSLT’14 DE-EN [28] is a German to English translation dataset that has 160k training examples. Our Transformer model is inherited from [3], which is a Post-LN Transformer placing its Layer Normalization after the summation of the skip connection and the residual branch. It has a $512$- dimensional word embedding layer and 1024 dimensions in its hidden FFN layer. It has $6$ encoder layers and $6$ decoder layers. We choose the learning rate to be $45e-$ 4 with inverse-sqrt learning schedule with $4000$ warmup updates and weight decay of $0.0001$.
Based on the discussion in the previous section, we have two ways of promoting emergence: A) promote emergence globally, by reducing the magnitude of weights in the encoder layers and increasing the magnitude of weights in the decoder layers; B) promote emergence clockwise, where we apply equation (8) to each encoder/ decoder block. In our transformer architecture, there is MLP block in each encoder/decoder layer consisting of $2$ layers. In our experiments, we can choose $\alpha$ up to $10$ for each MLP block, and notably we can see fast increase of BLEU score in the first few epochs. For example, with $\alpha=10$, the BLEU score after first epoch reaches $6.02$ while for T-Fixup the BLEU score after first epoch is $3.79$.
For global emergence promoting scheme, we first initialize our scheme based on T-Fixup, and then increase the magnitude of weights in the decoder layers by $2$. We run the models for the maximum of $80$ epochs and evaluate the BLEU score every epoch, and report the best BLEU scores throughout training for each run and the result is in Table 2.
5 Conclusion
In this paper, we introduced a novel and straightforward neural network initialization scheme inspired by the concept of emergence. Building on the emergent network measures proposed by Li (2023) [2], our method adjusts the layer-wise variance parameters to enhance the number of paths from inactive to active nodes, thereby achieving higher emergence values. This approach is not only easy to implement but also requires no additional optimization or training steps compared to conventional methods like Xavier and Kaiming initialization.
Our extensive evaluations across various architectures, including convolutional neural networks (CNNs) for image recognition and transformers for machine translation, demonstrate the significant advantages of our initialization scheme. The empirical results show that our method substantially improves model accuracy and convergence speed on standard datasets such as CIFAR-10 and the IWSLT-14 translation task.
By focusing on enhancing the emergent properties of neural networks from the outset, our initialization scheme facilitates better feature differentiation and integration, leading to more effective learning. The theoretical foundation and practical benefits of our method offer a promising direction for future research and applications in machine learning.
Our work contributes to the growing body of research that seeks to understand and harness the power of emergence in neural networks. By providing a simple yet powerful modification to existing initialization techniques, we open new avenues for improving neural network training methodologies. Future work could explore further optimizations and adaptations of our initialization scheme to other types of neural architectures and more complex tasks.
Overall, our emergence-promoting initialization scheme represents a valuable addition to current neural network initialization practices, offering both theoretical insights and practical improvements for the development of more robust and capable machine learning models.